OpenStudy (anonymous):

Let g be a differentiable function of one variable with g(2) = 2, g'(2) = -1. In what direction at the point (-1,-2) does the function f(x,y) =xg(y/x) increase most rapidly?

OpenStudy (anonymous):

In the direction of gradient

OpenStudy (anonymous):

take partial derivatic of f w.r.t x=g(y/x)+xg'(y/x)(-y/x^2)=g(y/x)-y/x g'(y/x) take partial derivative of w.r.t y=g'(y/x) so grad(f)=[g(y/x)+xg'(y/x)(-y/x^2)=g(y/x)-y/x g'(y/x)]i+[g'(y/x)]j at point (-1,-2) grad(f)=[g(2)-2g'(2)]i+g'(2)j =[2+2]i-j=4i-j